OpenStudy (karategirl2002):

The circle graph shows the number of gallons of ice cream ordered for the concert in the park. What percentage of the ice cream ordered was strawberry? A.40% B. 33 1/2% C. 9% D. 30% @Xmoses1 @esamalaa @jim_thompson5910 @RobertoTB @Squirrels

OpenStudy (karategirl2002):



OpenStudy (karategirl2002):

the graph

OpenStudy (karategirl2002):

@DeeDeeee_

OpenStudy (karategirl2002):

@minnie♥mouse

OpenStudy (karategirl2002):

@David.

OpenStudy (karategirl2002):

@RobertoTB

OpenStudy (anonymous):

the all =9+10+11=30 strawberry=9

OpenStudy (anonymous):

strawberry %=9/30 *100%